if sarah is two years old and her brother danayl is half her age, how old danyal be when when sarah turns 50 years old
Answer by henew(13)

Age of Sarah =2
Age of Danayl = half of 2 =1
When sarah turns 50, she will have lived another (in addition to her current age) 50-2= 48 years.
Then Danayl also will have lived another 48 years.
Age of Danayl will be 1+ 48= 49 years.

OR
Current age difference between Sarah and Danayl = 1 year.
This will remain constant throughout their life.( as both age equally)
So, when Sarah turns 50, Danayl will be 50 -1 = 49 years.